OpenStudy (anonymous):

Solve for x. 3^(x-2)=5^x

OpenStudy (anonymous):

\[3^{x-2}=5^x\]

OpenStudy (anonymous):

oh wait i think i got it

OpenStudy (anonymous):

\[\log_{3}5^x=x-2 \] \[x \log_{3}5=x-2 \] is this right?

OpenStudy (michele_laino):

we have: \[\left( \frac{ 3 }{ 5 } \right)^{x}=9\] then: \[x \log _{10}(3/5)=2 \log _{10}3\] and: \[x=\frac{ 2 \log _{10}3 }{ \log _{10}(3/5) }\]

OpenStudy (michele_laino):

yes! it is!

OpenStudy (anonymous):

Awesome, I totally forgot i had to use that rule